# Algebraic solution of project scheduling problems with temporal constraints

### v1

@@ Abstract
New solutions for problems in optimal scheduling of activities in a project under temporal constraints are developed in the framework of tropical algebra, which deals with the theory and application of algebraic systems with idempotent operations. We start with a constrained tropical optimization problem that has an objective function represented as a vector form given by an arbitrary matrix, and that can be solved analytically in a closed but somewhat complicated form. We examine a special case of the problem when the objective function is given by a matrix of unit rank, and show that the solution can be sufficiently refined in this case, which results in an essentially simplified analytical form and reduced computational complexity of the solution. We exploit the obtained result to find complete solutions of project scheduling problems to minimize the project makespan and the maximum absolute deviation of start times of activities under temporal constraints. The constraint under consideration include “start–start”, “start–finish” and “finish–start” precedence relations, release times, release deadlines and completion deadlines for activities. As an application, we consider optimal scheduling problems of a vaccination project in a medical centre.
@@ 1 Introduction
The scheduling problems, which have temporal objectives and constraints, and do not take into account other scarce resources (manpower, machines, tools, equipment, space, money), constitute an important class of temporal project scheduling problems [5, 21]. These problems may occur as auxiliary problems in the general project scheduling and are of independent interest.
@@
Since the early research works on tropical algebra in the 1960s [22, 4, 6], temporal scheduling problems have been applied to motivate and illustrate the development. In subsequent years, these problems continued to serve as important applications and to inspire further researches in the area (see, e.g., publications by [27, 2, 9, 24, 10]).
An efficient technique to handle temporal scheduling problems is proposed and developed in [13, 14, 16, 15, 17], based on the formulation and solution of the scheduling problems as optimization problems in the tropical algebra setting (tropical optimization problems). This technique offers direct solutions in an analytical vector form ready for further analysis with formal methods and numerical computation with moderate complexity, and thus can complement and supplement existing algorithmic approaches.
The analytical solution of the project scheduling problems may be rather cumbersome and involve many matrix-vector operations. However, as it is shown in [18] (see also some examples in [16, 15]), the solution of the problems with specific objectives, such as minimizing the absolute deviation of start times of activities, can be simplified.
@@
The solutions obtained extend and further develop the results in the paper [18]. In contrast to these results, we offer a solution to a tropical optimization problem that has a more general form of the objective function and an extended system of constraints. We apply this solution to new project scheduling problems with temporal constraints, which results in new complete solutions of the problems in a compact vector form.
The rest of the paper is organized as follows. In Section 2, we describe the temporal project scheduling problems of interest, and represent them in the usual algebraic form. In Section 3, we outline the key definitions, notations and results of tropical algebra to provide a formal framework for further discussion. Section 4 includes our main result, which provides a new solution to an optimization problem given in the tropical algebra setting. The result obtained is applied in Section 5 to derive solutions of the scheduling problems formulated above. In Section 6, we discuss the application of the solutions to the optimal scheduling of vaccination sessions and give illustrative numerical examples. Section 7 presents some conclusion remarks.
@@ 2 Project scheduling problems
We consider project scheduling problems with time-based objectives and constraints, which may typically occur in the project management practice in various areas [5, 21, 25, 26, 1].
@@
The scheduling problems under consideration are to determine the start and finish times for each activity to find optimal schedules that achieve specified objectives under the given temporal constraints. One of the objectives is set to minimize the project makespan, which constitutes a natural optimality criterion in many scheduling problems. Another objective is to minimize the absolute deviation of start times over all activities, and it is of interest when for some technological, organizational or other reasons, all activities in the project have to be scheduled to start as simultaneously as possible.
Consider a project that consists of $n$ activities and introduce the following notation. For each activity $i=1,\ldots,n$, we denote by $x_{i}$ the start time and by $y_{i}$ the finish time to be scheduled. Let $b_{i}$ be a given minimal allowed time lag between the start of activity $j$ and start of $i$, which specifies the “start–start” precedence relationship having the form of the inequality $b_{ij}+x_{j}\leq x_{i}$. If this lag is undefined, it is set to $\infty$- ∞. Combining such inequalities for all $j$ yields one equivalent inequality
@@
Each activity $i$ is assumed to finish immediately as all “start–finish” constraints are satisfied, which makes at least one inequality hold as an equality. As a result, we further replace the combined inequality by the equality
@@
Furthermore, given a minimal allowed time lag $d_{ij}$ between the finish of activity $j$ and start of $i$ ($c_{ij}=-\infty$ if undefined), the “finish–start” precedence relationship is the inequality $d_{ij}+y_{j}\leq x_{i}$, which yields the inequality
@@ 3.1 Idempotent semifield
Consider a set $\mathbb{X}$ that contains two distinct elements zero $\mathbb{0}$ and one $\mathbb{1}$, and is closed under addition $\oplus$ and multiplication $\otimes$. We assume that the triple $(\mathbb{X},\mathbb{0},\oplus)$ is a commutative idempotent monoid, $(\mathbb{X}\setminus\{\mathbb{0}\},\mathbb{1},\otimes)$ is an Abelian group, and multiplication is distributive over addition. The algebraic system $(\mathbb{X},\mathbb{0},\mathbb{1},\oplus,\otimes)$ is commonly referred to as the tropical (idempotent) semifield.
@@
Every nonzero $x$ has an inverse $x^{-1}$ such that $xx^{-1}=\mathbb{1}$. Taking inverse is antitone: for nonzero $x,y\in\mathbb{X}$, the inequality $x\leq y$ results in $x^{-1}\geq y^{-1}$. Integer exponents routinely represent iterative multiplication as $x^{p}=xx^{p-1}$, $x^{-p}=(x^{-1})^{p}$ and $\mathbb{0}^{p}=\mathbb{0}$ for any $x\neq\mathbb{0}$ and integer $p>0$. We assume that the equation $x^{p}=a$ has a unique solution for $x$ for any $a\in\mathbb{X}$ and integer $p>0$ to make the semifield radicable, where the rational exponents are defined as well.
For any $a,b\in\mathbb{X}$ and rational $q>0$, the binomial identity takes the form $(a\oplus b)^{q}=a^{q}\oplus b^{q}$, whereas the inequality between the arithmetic and geometric means becomes $a\oplus b\geq(ab)^{1/2}$. The extension of both identity and inequality to more than two summands is straightforward.
An example of a linearly ordered, radicable, idempotent semifield is the real system $\mathbb{R}_{\max,+}=(\mathbb{R}\cup\{-\infty\},-\infty,0,\max,+)$, also known as max-plus algebra. In this semifield, addition is defined as maximum and multiplication as arithmetic addition with the zero and one given respectively by $\infty$- ∞ and 00. The inverse for $x\neq\mathbb{0}$ corresponds to the conventional opposite number $-x$, the exponentiation $x^{y}$ coincides with the standard multiplication $xy$, and the order given by addition is equivalent to the natural linear order on $\mathbb{R}$.
@@ 3.2 Algebra of matrices and vectors
Consider matrices and vectors of elements from $\mathbb{X}$ and denote the set of matrices with $m$ rows and $n$ columns by $\mathbb{X}^{m\times n}$ and. Denote the set of column vectors with $n$ elements by $\mathbb{X}^{n}$. A matrix (vector) with all entries equal to zero $\mathbb{0}$ is the zero matrix (vector) denoted by $\bm{0}$. If a vector has no zero entries, it is called regular. A matrix without zero columns is called column-regular.
@@
For a vector $\bm{x}$ its transpose is denoted $\bm{x}^{T}$. The multiplicative conjugate transpose of a nonzero column vector $\bm{x}=(x_{i})$ is the row vector $\bm{x}^{-}=(x_{i}^{-})$, where $x_{i}^{-}=x_{i}^{-1}$ if $x_{i}\neq\mathbb{0}$, and $x_{i}^{-}=\mathbb{0}$ otherwise.
@@
A square matrix with all entries equal to $\mathbb{1}$ on the diagonal and to $\mathbb{0}$ elsewhere, is the identity matrix denoted $\bm{I}$. The nonnegative integer exponent of a matrix $\bm{A}\in\mathbb{X}^{n\times n}$ is defined for any integer $p>0$ as $\bm{A}^{p}=\bm{A}^{p-1}\bm{A}$, $\bm{A}^{0}=\bm{I}$.
@@ 4 Tropical optimization problems
In this section, we focus on optimization problems, which are formulated and solved in the tropical algebra setting, and hence referred to as tropical optimization problems. We start with a problem of constrained minimization of a multiplicative conjugate quadratic (pseudoquadratic) form $\bm{x}^{-}\bm{A}\bm{x}$, where $\bm{A}$ is a given square matrix and $\bm{x}$ is the unknown column vector. We describe an existing solution of the problem, known in a rather cumbersome matrix-vector form, and make some comments on the solution.
@@ 4.1 Minimization of conjugate quadratic forms
A complete solution of problem (10) can be obtained for instance as a direct consequence of the solution of a problem with an extended set of constraints or with a more general objective function in [14]. In what follows, we use the solution given in the form of the next statement.
@@
$\bm{x}=\bm{G}\bm{u},\qquad\bm{G}=\bigoplus_{k=1}^{n-1}\theta^{-k}\bm{G}_{k}\oplus\bm{B}^{\ast},$ (14)
@@
In the statement of the theorem, the condition $\mathop{\mathrm{missing}}{Tr}(\bm{B})\leq\mathbb{1}$ is responsible for the existence of regular solutions of the linear inequality $\bm{B}\bm{x}\leq\bm{x}$, whereas the condition $\bm{h}^{-}\bm{B}^{\ast}\bm{g}\leq\mathbb{1}$ ensures that both the linear inequality and the box-constraint $\bm{g}\leq\bm{x}\leq\bm{h}$ have a common regular solution.
Note that the solution provided by the theorem has a computational complexity of order not exceeding $O(n^{5})$ (see, e.g., [14, 15]).
@@ 4.2 Conjugate quadratic form with rank-one matrix
We start with the observation that under the condition $\bm{q}^{-}\bm{p}>\mathbb{0}$, the spectral radius of the matrix $\bm{p}\bm{q}^{-}$ is bounded from zero as
@@
After taking roots and summing over all $k=1,\ldots,n-1$, we represent the obtained sum as
@@
We use the last inequality to obtain an upper bound for the second terms as
@@
$\bm{G}=\theta^{-1}\bm{G}_{1}\oplus\bigoplus_{k=2}^{n-1}\theta^{-k}\bm{G}_{k}\oplus\bm{B}^{\ast},\qquad\bm{G}_{1}=\bigoplus_{0\leq i_{0}+i_{1}\leq n-2}\bm{B}^{i_{0}}\bm{p}\bm{q}^{-}\bm{B}^{i_{1}}.$
@@
To obtain the value of this sum, we calculate the vectors $\bm{B}^{i}\bm{p}$ and $\bm{B}^{j}\bm{g}$ by using matrix-vector multiplications consequently for all $i,j=0,\ldots,n-2$, which requires $O(n^{3})$ operations. With these vectors already computed, we find all summands and then the entire sum in $O(n^{2})$ time. Given the Kleene matrix $\bm{B}^{\ast}$, the second term of the form $\bm{q}^{-}\bm{B}^{\ast}\bm{p}$ requires $O(n^{2})$ operations. Considering the complexity of $O(n^{3})$ to obtain the matrix $\bm{B}^{\ast}$, we see that calculation with (17) takes not more than $O(n^{3})$ operations.
The most computationally demanding part of finding solution vectors by using (18) and (19) is calculating the matrix $\bm{G}$ and in particular the sum
$\bigoplus_{0\leq i+j\leq n-2}\theta^{-1}\bm{B}^{i}\bm{p}\bm{q}^{-}\bm{B}^{j}.$
After computing all vectors $\bm{B}^{i}\bm{p}$ and $\bm{q}^{-}\bm{B}^{j}$ in $O(n^{3})$ time, every summand in the above sum is calculated using $O(n^{2})$ operations. With the number of summands equal to $21+\cdots+n-1=n(n-1)/$ 2, the number of operations required to obtain the sum is estimated to be $O(n^{4})$. Since the complexity of calculating the matrix $\bm{B}^{\ast}$ is $O(n^{3})$, the complexity involved in evaluating the matrix $\bm{G}$, and hence solution vectors $\bm{x}$ is $O(n^{4})$. This complexity is higher than that required to evaluate the minimum $\theta$, and can be considered as the overall complexity of the solution provided by Theorem 3.

### v2

@@ Abstract
New solutions for problems in optimal scheduling of activities in a project under temporal constraints are developed in the framework of tropical algebra which deals with the theory and application of algebraic systems with idempotent operations. We start with a constrained tropical optimization problem that has an objective function represented as a vector form given by an arbitrary matrix, and that can be solved analytically in a closed but somewhat complicated form. We examine a special case of the problem when the objective function is given by a matrix of unit rank, and show that the solution can be sufficiently refined in this case, which results in an essentially simplified analytical form and reduced computational complexity of the solution. We exploit the obtained result to find complete solutions of project scheduling problems to minimize the project makespan and the maximum absolute deviation of start times of activities under temporal constraints. The constraints under consideration include “start–start”, “start–finish” and “finish–start” precedence relations, release times, release deadlines and completion deadlines for activities. As an application, we consider optimal scheduling problems of a vaccination project in a medical centre.
@@ 1 Introduction
Scheduling problems that have temporal objectives and constraints, and do not take into account other scarce resources (manpower, machines, tools, equipment, space, money), constitute an important class of temporal project scheduling problems [5, 22]. These problems may occur as auxiliary problems in the general project scheduling and are of independent interest.
@@
Since the early research works on tropical algebra in the 1960s [23, 4, 6], temporal scheduling problems have been applied to motivate and illustrate the development. In subsequent years, these problems continued to serve as important applications and to inspire further researches in the area (see, e.g., [28, 2, 9, 25, 10]).
An efficient technique to handle temporal scheduling problems is proposed and developed in [13, 14, 16, 15, 17], based on the formulation and solution of the scheduling problems as optimization problems in the tropical algebra setting (tropical optimization problems). The analytical solution of the project scheduling problems may be rather cumbersome and involve many matrix-vector operations. However, as it is shown in [19] (see also some examples in [16, 15]), the solution of the problems with specific objectives, such as minimizing the absolute deviation of start times of activities, can be simplified.
@@
The purpose of this study is to develop a general solution approach based on a tropical optimization technique proposed in [19] to handle a class of optimization problems that are frequently encountered in project scheduling. The technique uses a special form of the objective function in the problems, which allows one to simplify the formal representation and reduce the computational complexity of solutions. As an application example, a temporal scheduling problem to minimize the maximum absolute deviation of start times of activities is solved in [19] under few constraints on start and finish times of activities. To extend and generalize this result, we now offer an analytical solution to a tropical optimization problem with a general objective function and an extended system of constraints. We apply this solution to new temporal project scheduling problems (including a problem of common interest to minimize the project makespan), which results in new complete solutions of the problems in a compact vector form.
The contribution of the study is twofold. First, a new approach is developed to solve a range of temporal project scheduling problems analytically in an explicit form suitable for further analysis with formal methods and for numerical computations with moderate complexity. The proposed analytical solution has a strong potential to complement and supplement the existing approaches, which are based on numerical algorithms. Second, new solutions are obtained for a class of tropical optimization problems, which can find application to solve actual problems in operations research, management science and other fields. This result has implications for further development of methods and extension of applications of tropical optimization.
The rest of the paper is organized as follows. In Section 2, we describe the temporal project scheduling problems of interest and represent them in the usual algebraic form. In Section 3, we outline the key definitions, notations and results of tropical algebra to provide a formal framework for further discussion. Section 4 includes our main result, which provides a new solution to an optimization problem given in the tropical algebra setting. The result obtained is applied in Section 5 to derive solutions of the scheduling problems formulated above. In Section 6, we discuss the application of the solutions to the optimal scheduling of vaccination sessions and give illustrative numerical examples. Section 7 presents some conclusion remarks.
@@ 2 Project scheduling problems
We consider project scheduling problems with temporal (time-based) objectives and constraints, which may typically occur in the project management practice in various areas [5, 22, 26, 27, 1].
@@
The scheduling problems under consideration are to determine the start and finish times for each activity to find optimal schedules that achieve specified objectives under the given temporal constraints. One of the objectives is set to minimize the project makespan (total duration), which constitutes a natural optimality criterion in many scheduling problems. Another objective is to minimize the absolute deviation of start times over all activities, and it is of interest when for some technological, organizational or other reasons, all activities in the project have to be scheduled to start as simultaneously as possible.
Consider a project that involves $n$ activities, and introduce the following notation. For each activity $i=1,\ldots,n$, we denote by $x_{i}$ the start time and by $y_{i}$ the finish time to be scheduled. Let $b_{ij}$ be a given minimal allowed time lag between the start of activity $j$ and start of $i$, which specifies the “start–start” precedence relationship having the form of the inequality $b_{ij}+x_{j}\leq x_{i}$. If this lag is undefined, it is set to $\infty$- ∞. Combining such inequalities for all $j$ yields one equivalent inequality
@@
Each activity $i$ is assumed to finish immediately once all “start–finish” constraints are satisfied, which makes at least one inequality hold as an equality. As a result, we further replace the combined inequality by the equality
@@
Furthermore, given a minimal allowed time lag $d_{ij}$ between the finish of activity $j$ and start of $i$ ($d_{ij}=-\infty$ if undefined), the “finish–start” precedence relationship is the inequality $d_{ij}+y_{j}\leq x_{i}$, which yields the inequality
@@ 3.1 Idempotent semifield
Consider a set $\mathbb{X}$ that contains two distinct elements zero $\mathbb{0}$ and one $\mathbb{1}$, and is closed under addition $\oplus$ and multiplication $\otimes$. We assume that the triplet $(\mathbb{X},\mathbb{0},\oplus)$ is a commutative idempotent monoid, $(\mathbb{X}\setminus\{\mathbb{0}\},\mathbb{1},\otimes)$ is an Abelian group, and multiplication is distributive over addition. The algebraic system $(\mathbb{X},\mathbb{0},\mathbb{1},\oplus,\otimes)$ is commonly referred to as the tropical (idempotent) semifield.
@@
Every nonzero $x$ has an inverse $x^{-1}$ such that $xx^{-1}=\mathbb{1}$. Taking inverse is antitone: for nonzero $x,y\in\mathbb{X}$, the inequality $x\leq y$ results in $x^{-1}\geq y^{-1}$. Integer exponents routinely represent iterative multiplication as $x^{p}=xx^{p-1}$, $x^{-p}=(x^{-1})^{p}$ and $\mathbb{0}^{p}=\mathbb{0}$ for any $x\neq\mathbb{0}$ and integer $p>0$. We assume that the equation $x^{p}=a$ has a unique solution in $x$ for any $a\in\mathbb{X}$ and integer $p>0$ to make the semifield radicable, where the rational exponents are defined as well.
For any $a,b\in\mathbb{X}$ and rational $q>0$, the binomial identity takes the form $(a\oplus b)^{q}=a^{q}\oplus b^{q}$, whereas the inequality between the arithmetic and geometric means becomes $a\oplus b\geq(ab)^{1/2}$. The extensions of both identity and inequality to more than two summands are straightforward.
An example of a linearly ordered, radicable, idempotent semifield is the real system $\mathbb{R}_{\max,+}=(\mathbb{R}\cup\{-\infty\},-\infty,0,\max,+)$, also known as max-plus algebra. In this semifield, addition is defined as maximum and multiplication as arithmetic addition with the zero $\mathbb{0}$ and one $\mathbb{1}$ given respectively by $\infty$- ∞ and 00. The inverse for $x\neq\mathbb{0}$ corresponds to the conventional opposite number $-x$, the exponentiation $x^{y}$ coincides with the standard multiplication $xy$, and the order determined by addition is equivalent to the natural linear order on $\mathbb{R}$.
@@ 3.2 Algebra of matrices and vectors
Consider matrices and vectors of elements from $\mathbb{X}$ and denote the set of matrices with $m$ rows and $n$ columns by $\mathbb{X}^{m\times n}$, and the set of column vectors with $n$ elements by $\mathbb{X}^{n}$. A matrix (vector) with all entries equal to zero $\mathbb{0}$ is the zero matrix (vector) denoted by $\bm{0}$. If a vector has no zero entries, it is called regular. A matrix without zero columns is called column-regular.
@@
For a vector $\bm{x}$, its transpose is denoted $\bm{x}^{T}$. The multiplicative conjugate transpose of a nonzero column vector $\bm{x}=(x_{i})$ is the row vector $\bm{x}^{-}=(x_{i}^{-})$, where $x_{i}^{-}=x_{i}^{-1}$ if $x_{i}\neq\mathbb{0}$, and $x_{i}^{-}=\mathbb{0}$ otherwise.
@@
A square matrix with all entries equal to $\mathbb{1}$ on the diagonal and to $\mathbb{0}$ elsewhere, is the identity matrix denoted $\bm{I}$. The nonnegative integer exponent of a matrix $\bm{A}\in\mathbb{X}^{n\times n}$ is defined for any integer $p>0$ as $\bm{A}^{p}=\bm{A}^{p-1}\bm{A}$ and $\bm{A}^{0}=\bm{I}$.
@@ 4 Tropical optimization problems
In this section, we focus on optimization problems, which are formulated and solved in the tropical algebra setting, and hence referred to as tropical optimization problems. We start with a problem of constrained minimization of a multiplicative conjugate quadratic (pseudoquadratic) form $\bm{x}^{-}\bm{A}\bm{x}$, where $\bm{A}$ is a given square matrix and $\bm{x}$ is an unknown column vector. We describe an existing solution of the problem, known in a rather cumbersome matrix-vector form, and make some comments on the solution.
@@ 4.1 Minimization of conjugate quadratic forms
A complete solution of problem (10) can be obtained for instance as a direct consequence of the solution of a problem with an extended set of constraints or with a more general objective function in [14]. In what follows, we use the solution given in the next claim.
@@
$\bm{x}=\bm{G}\bm{u},\qquad\bm{G}=(\theta^{-1}\bm{A}\oplus\bm{B})^{\ast}=\bigoplus_{k=1}^{n-1}\theta^{-k}\bm{G}_{k}\oplus\bm{B}^{\ast},$ (14)
@@
In the statement of the theorem, the condition $\mathop{\mathrm{missing}}{Tr}(\bm{B})\leq\mathbb{1}$ is responsible for the existence of regular solutions of the linear inequality $\bm{B}\bm{x}\leq\bm{x}$, whereas the condition $\bm{h}^{-}\bm{B}^{\ast}\bm{g}\leq\mathbb{1}$ ensures that both the linear inequality constraint and the box constraint $\bm{g}\leq\bm{x}\leq\bm{h}$ have a common regular solution.
Note that the solution provided by the theorem has a computational complexity of order not exceeding $O(n^{5})$, where the most computationally expensive part which yields the upper bound is the evaluation of the minimum $\theta$ (see, e.g., [14, 15, 18]).
@@ 4.2 Conjugate quadratic form with rank-one matrix
We start with the observation that under the condition $\bm{q}^{-}\bm{p}>\mathbb{0}$, the spectral radius of the matrix $\bm{p}\bm{q}^{-}$ is greater than zero as
@@
After taking roots and summing over all $k=1,\ldots,n-1$, we arrive at the sum
@@
We use the last inequality to derive an upper bound for the second terms as
@@
$\bm{G}=(\theta^{-1}\bm{p}\bm{q}^{-}\oplus\bm{B})^{\ast}=\theta^{-1}\bm{G}_{1}\oplus\bigoplus_{k=2}^{n-1}\theta^{-k}\bm{G}_{k}\oplus\bm{B}^{\ast},$
where
$\bm{G}_{1}=\bigoplus_{0\leq i_{0}+i_{1}\leq n-2}\bm{B}^{i_{0}}\bm{p}\bm{q}^{-}\bm{B}^{i_{1}}.$
@@
To obtain the value of this sum, we calculate the vectors $\bm{B}^{i}\bm{p}$ and $\bm{B}^{j}\bm{g}$ by using matrix-vector multiplications consequently for all $i,j=0,\ldots,n-2$, which requires $O(n^{3})$ operations. With these vectors already computed, we find all summands and then the entire sum in $O(n^{2})$ time. Given the Kleene matrix $\bm{B}^{\ast}$, the second term of the form $\bm{q}^{-}\bm{B}^{\ast}\bm{p}$ requires $O(n^{2})$ operations. Considering the complexity of $O(n^{3})$ to obtain the matrix $\bm{B}^{\ast}$, we see that calculation with (17) takes no more than $O(n^{3})$ operations.
The most computationally demanding part of finding solution vectors by using (18) and (19) is calculating the matrix $\bm{G}$. It follows from the proof of Theorem 3 that this matrix is initially defined as the Kleene matrix $\bm{G}=(\theta^{-1}\bm{p}\bm{q}^{-}\oplus\bm{B})^{\ast}$. With the minimum $\theta$ fixed, the complexity of calculating this matrix is no more than $O(n^{3})$. This complexity has the same order as that required to evaluate the minimum $\theta$, and can thus be considered as the overall complexity of the solution provided by Theorem 3.
Finally, we note that in the context of project scheduling problems, the constraint matrix $\bm{B}$ frequently appears to be sparse (have few nonzero entries), which allows one to achieve further reduction of computational complexity by using efficient algebraic techniques specifically designed for sparse matrices.